OpenStudy (anonymous):

Solve using substitution: 4x - y - 4 = 0 7y = 29 + 9x

OpenStudy (anonymous):

pls I need help

OpenStudy (anonymous):

Infinitely many solutions No solution (3, 8) (5, 16)

OpenStudy (anonymous):

y=4x-4 substituting in the second equation,,we get 7(4x-4) =29+9x expanding it,, 28x-28=29+9x 19x=57 x=3 replacing x in the first equation,, y =8 answer = 3,8